# Review-Based Cross-Domain Recommendation via Hyperbolic Embedding and Hierarchy-Aware Domain Disentanglement

### v1

@@ Abstract.
The issue of data sparsity poses a significant challenge to recommender systems. In response to this, algorithms that leverage side information such as review texts have been proposed. Furthermore, Cross-Domain Recommendation (CDR), which captures domain-shareable knowledge and transfers it from a richer domain (source) to a sparser one (target), has received notable attention. Nevertheless, the majority of existing methodologies assume a Euclidean embedding space, encountering difficulties in accurately representing richer text information and managing complex interactions between users and items. This paper advocates a hyperbolic CDR approach based on review texts for modeling user-item relationships. We first emphasize that conventional distance-based domain alignment techniques may cause problems because small modifications in hyperbolic geometry result in magnified perturbations, ultimately leading to the collapse of hierarchical structures. To address this challenge, we propose hierarchy-aware embedding and domain alignment schemes that adjust the scale to extract domain-shareable information without disrupting structural forms. The process involves the initial embedding of review texts in hyperbolic space, followed by feature extraction incorporating degree-based normalization and structure alignment. We conducted extensive experiments to substantiate the efficiency, robustness, and scalability of our proposed model in comparison to state-of-the-art baselines.
@@
copyright: acmcopyrightjournalyear: 2024conference: Proceedings of the 33rd ACM International Conference; October 21–25, 2024; Boise, Idaho, USAbooktitle: Proceedings of the 33rd ACM International Conference ’24, October 21–25, 2024, Boise, Idaho, USAprice: 15.00isbn: 978-1-4503-9236-5/22/10doi: 10.1145/3511808.3557324ccs: Information systems Recommender systems
@@ 1. Introduction
A recommender system has evolved into a fundamental tool across real-world applications (Lu et al., 2015; Zhang et al., 2019), exemplified by its integration into platforms such as Amazon, and Tripadvisor. Despite remarkable popularity and commercial successes, the performance can be inflicted heavily in data-scarce scenarios. Weathering of the data sparsity impediments, encompassing issues like cold-start problems, has emerged as a central focus of scholarly inquiry. Numerous research endeavors have employed various forms of side information ranging from social relationships (Kazienko et al., 2011), and hierarchical interactions (Liu et al., 2019) to item images (Addagarla and Amalanathan, 2020) aiming to augment recommendation accuracy. In particular, textual data in the form of reviews stands out as one of the most extensively utilized side information. The realm of text-based recommender systems has witnessed significant advancements as evidenced by investigations such as (Zheng et al., 2017), (Al-Ghuribi and Noah, 2019), (Srifi et al., 2020), and (Choi et al., 2022). These endeavors have demonstrated a measure of effectiveness in ameliorating sparsity concerns. However, inherent limitations persist in addressing fundamental issues, especially when the extent of interaction within a dataset is insufficient.
Figure 1. The geometric properties of Euclidean ($\mathbb{E}$, left) and Hyperbolic spaces ($\mathbb{H}$, right). A hyperbolic space leverages the advantages of a wide space by placing nodes with high degrees close to the origin. However, common methods in recommender systems bring the relevant nodes closer, leading to a structural collapse in the hyperbolic geometry
@@
To handle this problem, recent method strides in Cross-Domain Recommendation (CDR) (Hu et al., 2018; Yuan et al., 2019; Fu et al., 2019) and transfer algorithms (Yuan et al., 2020; Yang et al., 2021; Hande et al., 2021). These methodologies commonly exploit the information from a source domain, characterized by abundant interactions relative to a target domain, to extract domain-shareable information. Some approaches concentrate on duplicate users across both domains (Yuan et al., 2019; Li and Tuzhilin, 2020; Liu et al., 2021; Guo et al., 2021). However, it is essential to note that these user binding strategies may face constraints arising from the absence of overlapping (duplicate) users and domain selection restrictions (Kang et al., 2019). Alternatively, more flexible methods that operate independent of specific users or contexts have been introduced (Cai et al., 2019; Zhao et al., 2020; Krishnan et al., 2020). This trend has prompted the development of disentangled representation learning techniques (Gretton et al., 2005; Bousmalis et al., 2016; Li et al., 2019; Peng et al., 2019), which can concurrently extract both domain-specific and domain-shareable knowledge. These algorithms empower the distillation of meaningful insights from seemingly counterproductive domains. More recently, novel approaches contrived on review-based disentangled representation learning (Cao et al., 2022; Choi et al., 2022), free of duplicate users or contexts, have demonstrated state-of-the-art performance.
We have discussed how recent studies mitigate the data sparsity problem by leveraging side information and employing domain disentanglement for knowledge transfer. While these methods are effective, a drawback exists as they rely on Euclidean geometry for the embedding (Globerson et al., 2004; Khoshneshin and Street, 2010), which has shown to be inadequate for modeling the user-item bipartite graphs whose node degrees follow a power-law distribution (Clauset et al., 2009). This inadequacy leads to a distortion in embedding due to the exponential growth in the graph’s volume with its radius. To solve this limitation, recent studies propose the adoption of hyperbolic geometry (Gulcehre et al., 2018; Khrulkov et al., 2020; Feng et al., 2020). In Figure 1, we present illustrative examples of Euclidean space ($\mathbb{E}$, left) and hyperbolic spaces ($\mathbb{H}$, right). As previously mentioned, the volume of hyperbolic space (distance from the origin) exhibits exponential growth ($e^{x}$) compared to the polynomial growth ($x$) in Euclidean space as they expand towards the boundary (Nickel and Kiela, 2017).
@@
Considering these factors, we focus on hyperbolic CDR, a relatively underexplored area (Xu and Cai, 2023). Despite the existence of various CDR methods, the majority of them bring features extracted from two closely related domains (Cao et al., 2022; Zhao et al., 2023), potentially provoking the loss of hierarchical information. As shown in the right side of Figure 1 ($\mathbb{H}$), directly reducing the distance between two items $i_{1}$ and $i_{2}$, purchased by the same user without considering their positions can detrimentally impact the entire representation. In detail, given that the degree of node $i_{1}$ is greater than node $i_{2}$, placing a node $i_{2}$ with fewer interactions closer to the origin than $i_{1}$ results in the loss of the advantages of an exponentially increasing space. In this context, there arises a need to devise clever mechanisms that facilitate knowledge transfer while preserving a tree-like structure. For this, we propose a solution equipped with two strategies; degree-based normalization and structure alignment. The proposed scheme not only preserves the position but also enhances the performance of domain disentanglement. Our approach is substantiated through theoretical understanding and empirical evidence from various experiments. In summary, our contributions are outlined below.
@@
  • We propose degree-based hierarchy alignment and scale adjustment to enhance the knowledge transfer between two domains. To our knowledge, this is the first attempt that achieves domain disentanglement in a hyperbolic space.
  • We present theoretical understandings that emphasize the importance of hierarchy preservation and domain disentanglement for recommender systems.
  • We conducted various experiments to verify the effectiveness of our method and the correctness of the theoretical analysis.
@@ 2. Related Work
In this section, we introduce the advent of recommender algorithms categorizing them as follows; (1) Review-based recommendations that utilize review texts as side information, (2) Cross-domain recommendations that exploit additional information from source domains for knowledge transfer, and (3) Hyperbolic recommendation which project the user-item representations onto hyperbolic space instead of Euclidean space to preserve hierarchical information.
@@ 2.1. Review-Based Recommendation
The explosive progress in text convolution techniques has ignited great attention in review-based recommender systems (Zheng et al., 2017; Chen et al., 2019b, a; Dong et al., 2020). For example, DeepCoNN (Zheng et al., 2017) employs two parallel convolutional neural networks (CNNs), and others further utilize an attention mechanism to exploit important words (Seo et al., 2017; Chen et al., 2018; Tay et al., 2018; Dong et al., 2020). Although these methods highlight the importance of review texts, their major limitation lies in the confined scale of target domains and the transfer of noisy data (Sachdeva and McAuley, 2020; Zeng et al., 2021). To address these issues, cross-domain recommendation and disentanglement techniques have emerged to acquire useful information from domains with richer interactions.
@@ 2.2. Cross-Domain Recommendation
Cross-domain recommendation (CDR) involves the utilization of additional information in other domains called source domains to address the sparseness in a target domain. Generally, these methods capture latent information from rating matrices or review texts (Fu et al., 2019; Krishnan et al., 2020) to capture and transfer knowledge authored by overlapping users in both source and target domains (Elkahky et al., 2015; Man et al., 2017; Zhu et al., 2021). Though certain methods (Wang et al., 2018; Zhao et al., 2020) underscore the significance of employing non-overlapping users for generalization, they lack examination of what information would be efficiently transferred. Thus, recent studies have shifted focus towards identifying the most relevant aspects between two domains, commonly referred to as domain-shareable features. The foundational mechanism often commences with domain adaptation (Ramakrishnan et al., 2018; Yuan et al., 2019; Bonab et al., 2021; Chen et al., 2023), which captures domain-shareable features through adversarial training. Advanced techniques that extract domain-specific and domain-shareable features simultaneously have been introduced more recently, collectively known as disentangled representation learning. These include MMT (Krishnan et al., 2020), DADA (Peng et al., 2019), DisenCDR (Cao et al., 2022), and SER (Choi et al., 2022). The pivotal aspect of this approach lies in domain disentanglement which strives to identify useful information for knowledge transfer.
@@ 2.3. Hyperbolic Recommendation
Most prior research on recommendation systems has primarily been conducted in a Euclidean space. These methods have demonstrated decent performance, but a challenge has been raised regarding their adequacy in modeling hierarchical structures such as user-item interactions or word vectors (Tifrea et al., 2018). Recent studies (Feng et al., 2020; Wang et al., 2021; Su et al., 2023) utilize a hyperbolic space for representation learning, including informative collaborative filtering (Yang et al., 2022a; Li et al., 2022; Wang et al., 2023) with geometric regularization (Yang et al., 2022b). However, most of them utilize a single-domain dataset, which is susceptible to data sparsity. To address this problem, several researchers have integrated CDR with hyperbolic space embedding (Xu and Cai, 2023; Guo et al., 2023), but they disregard domain disentanglement and hierarchy structure preservation, which are particularly crucial for accurate knowledge transfer in an exponentially expanded space. Thus, we aim to suggest a new hyperbolic CDR that preserves the structural property to better embed the user-item interactions.
@@ 3. Preliminaries
We first introduce the basic concepts of manifolds and hyperbolic geometry. Differential geometry defines three space types, hyperbolic, Euclidean, and spherical, based on curvatures. Especially, a hyperbolic space is one type of non-Euclidean space, which has a constant negative curvature at all points. In the literature, various mathematical formulations can be utilized to describe hyperbolic spaces, such as the Riemannian manifold (Lin and Zha, 2008), Poincaré ball (Nickel and Kiela, 2017), and Lorentz model (Nickel and Kiela, 2018). Here, we employ the Poincaré ball for the ease of visualization (Gulcehre et al., 2018) and the Lorentz model for numerical operation (Law et al., 2019), respectively.
@@
where $k$ is the radius of the ball. The above equation converts a Euclidean metric tensor $g^{E}$ to hyperbolic one. If $k=0$, we can easily infer that the ball is identical to the Euclidean space. Also, a distance function on $\mathcal{P}$ is defined as below:
@@
Based on these equations, we will explain methods for knowledge transfer in hyperbolic space. In Table 1, we describe the notations that will be used throughout this paper.
Table 1. Notations
Symbol Explanation
$\mathbb{E},\mathbb{H}$ Euclidean and Hyperbolic spaces
$\mathrm{u,i,y_{u,i},r_{u,i}}$ User, item, rating, and individual review
$\mathrm{R}$ Aggregated reviews (documents)
$\mathrm{S,T}$ Domain-specific features extracted from $\mathrm{R}$
$\mathrm{\widehat{S},\widehat{T}}$ Domain-shareable features extracted from $\mathrm{R}$
$\mathrm{N_{s},N_{t}}$ Batch of source and target domain datasets
$\mathrm{exp_{x}(v)}$ Mapping from Euclidean to hyperbolic space
$\mathrm{log_{x}(v)}$ Mapping from hyperbolic to Euclidean space
$\mathrm{d_{\mathcal{P}}(x,y)}$ Distance measure in Poincaré ball
$\mathrm{d_{\mathcal{L}}(x,y)}$ Distance measure in Lorentz model
$\mathrm{\mathcal{L}}$ Loss function
$\mathrm{\lambda}$ Hyperparameters in $\mathcal{L}$
@@ 4. Methodology
  • Word embedding. This part vectorizes reviews using pre-trained word embedding. In contrast to the previous methods that adopt Euclidean embedding such as word2vec [^1] https://code.google.com/archive/p/word2vec (Mikolov et al., 2013) or GloVe [^2] https://nlp.stanford.edu/projects/glove (Pennington et al., 2014), we employ the Poincaré Glove [^3] https://github.com/alex-tifrea/poincare_glove (Tifrea et al., 2018) which better preserves the hierarchical property of words.
  • Feature extraction. This module elicits pertinent information from embedded documents using three types of feature extractors (FEs); the shared FE focuses on domain-shareable knowledge for transfer while the source and target FEs capture the domain-specific features.
@@ 4.1. Word Embedding
Each data record follows a format (u, i, $y_{u,i}$, $r_{u,i}$), which means that a user $u$ purchased an item $i$ and left a rating $y_{u,i}$ and a review $r_{u,i}$. Assume that we are concerned with a possibly unseen rating $y_{u,i}$. We first aggregate all reviews of $u$ and $i$. Specifically, for a user $u$, we gather all reviews written by her except for the specific pair $r_{u,i}$ (not available during inference) and consider them as a single document $R_{u}$. Likewise, one can construct the collection of item reviews, $R_{i}$. Here, we ignore the temporal sequence of ratings or reviews. Finally, we apply the word embedding function to $R_{u}$ and $R_{i}$. Although many strategies are applicable (e.g, word2vec [^4] https://code.google.com/archive/p/word2vec (Mikolov et al., 2013)), we focus on the Euclidean (Pennington et al., 2014) and Poincaré Glove (Tifrea et al., 2018) here.
@@
    where $b$ is the bias.
@@
    where $d_{\mathcal{P}}$ is a distance function in Eq. 2 and $h(x)=cosh^{2}$.
The Euclidean and Poincaré renditions  [^5] https://polybox.ethz.ch/index.php/s/TzX6cXGqCX5KvAn have been trained using 1.4 billion tokens from English Wikipedia, and we employ them as pre-trained word embedding $f(\cdot)$ (please see Table 4.5). Using this, the textual documents $R_{u}$ and $R_{i}$ are mapped into the matrix $\mathrm{R^{\mathbb{E}}_{u}=f(R_{u}),R^{\mathbb{E}}_{i}=f(R_{i})}$ of $\mathcal{R}^{n\times d}$, where $n$ is the vocabulary size and $d$ is the embedding dimension. We project these matrices onto the hyperbolic space using the exponential map in Eq. 6 as below:

### v3

# Review-Based Hyperbolic Cross-Domain Recommendation

@@ Abstract.
The issue of data sparsity poses a significant challenge to recommender systems. Recently, algorithms that leverage side information (review texts) or Cross-Domain Recommendation (CDR) have emerged. Nevertheless, existing methodologies assume an Euclidean embedding space, encountering difficulties in accurately representing richer text information and managing complex user-item interactions. This paper advocates a hyperbolic CDR approach for modeling review-based user-item relationships. We first emphasize that conventional distance-based domain alignment techniques may cause problems because small modifications in hyperbolic geometry result in magnified perturbations, ultimately leading to the collapse of hierarchical structures. To address this challenge, we propose hierarchy-aware embedding and domain alignment schemes that adjust the scale to extract domain-shareable information without disrupting structural forms. Extensive experiments substantiate the efficiency, robustness, and scalability of the proposed model. The source code is given here [^1] https://github.com/ChoiYoonHyuk/HEAD.
@@
ccs: Information systems Recommender systems
Figure 1. The geometric properties of hyperbolic space require that popular (or most interacted) nodes be placed near the origin. Let us assume that user $u_{1}$ purchased an item $i_{1}$ (left). Given that general algorithms bring relevant nodes closer together, after the update, a structural collapse occurs (right) as $i_{1}$ moves farther from the origin
@@ 1. Introduction
A recommender system has evolved into a fundamental tool across real-world applications (Lu et al., 2015; Zhang et al., 2019) such as Amazon, and Tripadvisor. Despite remarkable popularity and commercial successes, the performance can be inflicted heavily in data-scarce scenarios. The impediments of data sparsity encompass issues like cold-start problems, which have gained academic focus recently. Numerous research endeavors have employed various forms of side information ranging from social relationships (Kazienko et al., 2011), and hierarchical interactions (Liu et al., 2019) to item images (Addagarla and Amalanathan, 2020). In particular, textual data in the form of reviews has become one of the most extensively utilized sources (Zheng et al., 2017; Al-Ghuribi and Noah, 2019; Srifi et al., 2020; Choi et al., 2022). These endeavors have demonstrated a measure of effectiveness in ameliorating sparsity concerns. However, inherent limitations persist in addressing fundamental issues, especially when the extent of interaction is insufficient.
@@
To handle this problem, recent method strides in Cross-Domain Recommendation (CDR) (Hu et al., 2018; Yuan et al., 2019; Fu et al., 2019; Yuan et al., 2020; Yang et al., 2021; Hande et al., 2021). These algorithms commonly exploit the information from a source domain, characterized by abundant interactions relative to a target domain to extract domain-shareable information. Some approaches concentrate on duplicate users across both domains (Yuan et al., 2019; Li and Tuzhilin, 2020; Liu et al., 2021; Guo et al., 2021). However, it is essential to note that these user-binding strategies may face constraints arising from the absence of overlapping (duplicate) users (Kang et al., 2019). Alternatively, more flexible methods that operate independent of specific users or contexts have been introduced (Cai et al., 2019; Zhao et al., 2020; Krishnan et al., 2020). This trend has prompted the development of disentangled representation learning techniques (Gretton et al., 2005; Bousmalis et al., 2016; Li et al., 2019; Peng et al., 2019), which can concurrently extract both domain-specific and domain-shareable knowledge. More recently, novel approaches contrived on review-based disentangled representation learning (Cao et al., 2022a; Choi et al., 2022) free of duplicate users or contexts have demonstrated state-of-the-art performance.
Nonetheless, the above methods rely on Euclidean geometry for embedding (Globerson et al., 2004; Khoshneshin and Street, 2010), which has been shown to be inadequate for modeling user-item bipartite graphs whose node degrees follow a power-law distribution (Clauset et al., 2009). This inadequacy leads to distortions in the embedding due to the exponential growth in the graph’s volume with its radius. In addition, the application of hyperbolic geometry to cross-domain recommendation (CDR) remains relatively unexplored (Xu and Cai, 2023), whereas recent studies on hyperbolic recommendation (Gulcehre et al., 2018; Khrulkov et al., 2020; Feng et al., 2020; Vinh Tran et al., 2020) focus on a single domain.
@@
In this paper, we propose a novel strategy for hyperbolic CDR since previous methods rely on features extracted from two closely related domains (Cao et al., 2022a; Zhao et al., 2023b, a), potentially losing hierarchical information. For example, directly reducing the distance between two items, $i_{1}$ and $u_{1}$ (Figure 1), purchased by a user without considering their positions can negatively impact the overall representation. As illustrated, distance-based minimization can lead to hierarchical collapse, as node positions cannot be preserved in this process. Specifically, if the degree of node $i_{1}$ is greater than others, placing a node with fewer interactions $u_{1}$ closer to the origin decreases the advantages of an exponentially increasing space. This context highlights the need for clever mechanisms that facilitate knowledge transfer while preserving a tree-like structure. To address this, we propose a solution with two strategies: degree-based normalization and structure alignment, substantiated through theoretical insights and empirical evidence from various experiments. In summary, our contributions are outlined as follows:
@@
  • We propose degree-based hierarchy alignment and scale adjustment to enhance the knowledge transfer between two domains. To our knowledge, this is the first attempt to achieve domain disentanglement in a hyperbolic space.
  • We present theoretical understandings to prove the importance of hierarchy preservation in domain disentanglement.
  • We conducted various experiments to verify the effectiveness of our method and the accuracy of the theoretical analysis.
@@ 2. Related Work
In this section, we introduce the advent of recommender algorithms categorizing them as follows; (1) Review-based recommendations (using side information), (2) Cross-domain recommendations (knowledge transfer), and (3) Hyperbolic recommendations that can preserve hierarchical information.
@@ 2.1. Review-Based Recommendation
The explosive progress in text convolution techniques has ignited great attention in review-based recommender systems (Zheng et al., 2017; Chen et al., 2019b, a; Dong et al., 2020). For example, DeepCoNN (Zheng et al., 2017) employs two parallel convolutional neural networks (CNNs), and others further utilize an attention mechanism to exploit important words (Seo et al., 2017; Chen et al., 2018; Tay et al., 2018; Dong et al., 2020). Although these methods highlight the importance of review texts, their major limitation lies in the confined scale of target domains and the transfer of noisy data (Sachdeva and McAuley, 2020; Zeng et al., 2021). To address these issues, cross-domain recommendation and disentanglement techniques have emerged to acquire useful information from richer domains.
@@ 2.2. Cross-Domain Recommendation
Cross-domain recommendation (CDR) utilizes additional information in extra (source) domain to address the sparseness in a target. Generally, these methods capture latent information from rating matrices or review texts (Fu et al., 2019; Krishnan et al., 2020) to capture and transfer knowledge authored by overlapping users in both source and target domains (Elkahky et al., 2015; Man et al., 2017; Zhu et al., 2021). Though certain methods (Wang et al., 2018; Zhao et al., 2020) underscore the significance of employing non-overlapping users for generalization, they lack examination of what information would be efficiently transferred. Thus, recent studies have shifted focus towards identifying the most relevant aspects between two domains, commonly referred to as domain-shareable features. The foundational mechanism often commences with domain adaptation (Ramakrishnan et al., 2018; Yuan et al., 2019; Bonab et al., 2021; Cao et al., 2022b; Chen et al., 2023), which captures domain-shareable features through adversarial training. Advanced techniques that extract domain-specific and domain-shareable features simultaneously have been introduced more recently, collectively known as disentangled representation learning. These include MMT (Krishnan et al., 2020), DADA (Peng et al., 2019), DisenCDR (Cao et al., 2022a), and SER (Choi et al., 2022). The pivotal aspect of this approach lies in domain disentanglement which strives to identify useful information for knowledge transfer.
@@ 2.3. Hyperbolic Recommendation
Most prior research on recommendation systems has primarily been conducted in an Euclidean space. These methods have demonstrated decent performance, but a challenge has been raised regarding their adequacy in modeling hierarchical structures such as user-item interactions or word vectors (Tifrea et al., 2018). Recent studies (Feng et al., 2020; Wang et al., 2021; Su et al., 2023) utilize a hyperbolic space for representation learning, including informative collaborative filtering (Yang et al., 2022a; Li et al., 2022; Wang et al., 2023) with geometric regularization (Yang et al., 2022b). However, most of them utilize a single-domain dataset, which is susceptible to data sparsity. To address this problem, several researchers have integrated CDR with hyperbolic space embedding, but they disregard domain disentanglement (Xu and Cai, 2023; Guo et al., 2023) and hierarchy structure preservation (Zhang et al., 2022), which are particularly crucial for accurate knowledge transfer in an exponentially expanded space. Thus, we aim to suggest a new hyperbolic CDR that preserves the structural property to better embed the user-item interactions.
@@ 3. Preliminaries
We start with the basic concepts of manifolds and hyperbolic geometry. Differential geometry defines three space types: hyperbolic, Euclidean, and spherical, based on curvatures. Especially, a hyperbolic space is one type of non-Euclidean space, which has a constant negative curvature at all points. In the literature, various mathematical formulations can be utilized to describe hyperbolic spaces, such as the Riemannian manifold (Lin and Zha, 2008), Poincaré ball (Nickel and Kiela, 2017), and Lorentz model (Nickel and Kiela, 2018). We employ the Poincaré ball for visualization (Gulcehre et al., 2018) and the Lorentz model for numerical operation (Law et al., 2019), respectively.
@@
where $k$ is the radius of the ball. The above equation converts an Euclidean metric tensor $g^{E}$ to a hyperbolic one. If $k=0$, we can easily infer that the ball is identical to the Euclidean space. Also, a distance function on $\mathcal{P}$ is defined as below:
@@ 4. Methodology
  • Word embedding. This part vectorizes reviews using pre-trained word embedding. In contrast to the previous methods that adopt Euclidean embedding such as word2vec [^2] https://code.google.com/archive/p/word2vec (Mikolov et al., 2013) or Euclidean GloVe [^3] https://nlp.stanford.edu/projects/glove (Pennington et al., 2014), we employ the Poincaré Glove [^4] https://github.com/alex-tifrea/poincare_glove (Tifrea et al., 2018) which better preserves the hierarchical property of words.
  • Feature extraction. This module elicits pertinent information from embedded documents using three types of feature extractors (FEs) (Choi et al., 2022); the shared FE focuses on domain-shareable knowledge for transfer while the source and target FEs capture the domain-specific features.
@@ 4.1. Word Embedding
Each data record follows a format (u, i, $y_{u,i}$, $r_{u,i}$), which means that a user $u$ purchased an item $i$ and left a rating $y_{u,i}$ and a review $r_{u,i}$. Assume that we are concerned with a possibly unseen rating $y_{u,i}$. We first aggregate all reviews of $u$ and $i$. Specifically, for a user $u$, we gather all reviews written by her except for the specific pair $r_{u,i}$ (not available during inference) and consider them as a single document $R_{u}$. Likewise, one can construct the collection of item reviews, $R_{i}$. Here, we ignore the temporal sequence of ratings or reviews. Finally, we apply the word embedding function to $R_{u}$ and $R_{i}$. Although many strategies are applicable (e.g, word2vec [^5] https://code.google.com/archive/p/word2vec (Mikolov et al., 2013)), we focus on the Euclidean- (Pennington et al., 2014) and Poincaré Glove (Tifrea et al., 2018) here.
@@
    where $b$ stands for the bias term.
@@
    where $d_{\mathcal{P}}$ is a distance function in Eq. 2 and $h(x)=cosh^{2}(x)$.
The Euclidean and Poincaré renditions  [^6] https://polybox.ethz.ch/index.php/s/TzX6cXGqCX5KvAn have been trained using 1.4 billion tokens from English Wikipedia, and we compare them as pre-trained word embedding $f(\cdot)$ in Table 5. Using this, the textual documents $R_{u}$ and $R_{i}$ are mapped into the matrix $\mathrm{R^{\mathbb{E}}_{u}=f(R_{u}),R^{\mathbb{E}}_{i}=f(R_{i})}$ of $\mathcal{R}^{n\times d}$, where $n$ is the vocabulary size and $d$ is the embedding dimension. We project these matrices onto the hyperbolic space using the exponential map in Eq. 6 as below: